OpenStudy (asylum15):

Using the laws of indicies, simplify, giving your answers with positive indices:

OpenStudy (asylum15):

\[18^{x} \times 9^{2+x}\over3^{3x-1} \times 6^{x+1}\]

OpenStudy (asylum15):

Hi cwr :)

OpenStudy (cwrw238):

hi

OpenStudy (asylum15):

Tricky problem?

OpenStudy (cwrw238):

yea tis a bit

OpenStudy (cwrw238):

we can write 9^2+x as 3 ^(4+2x)

OpenStudy (cwrw238):

that gives us 3^(4+ 2x - (3x-1) = 3^(5-x)

OpenStudy (cwrw238):

= 18^x * 3^(5-x) ------------ 6^(x+1)

OpenStudy (asylum15):

Thats it I guess? :)

OpenStudy (cwrw238):

well - its a bit more simple - lol